OpenStudy (anonymous):

Simple math question Please help thanks... Chang has 2 shirts: a white one and a black one. He also has 2 pairs of pants, one blue and one tan. What is the probability, if Chang gets dressed in the dark, that he winds up wearing the white shirt and tan pants? Show your work.

OpenStudy (anonymous):

@DanJS @kitty_lover1234

OpenStudy (danjs):

probability of the first thing times probability of the second occuring

OpenStudy (anonymous):

i dont understand?

OpenStudy (danjs):

he has 2 shirts, each one has a 1 in 2 chance of being picked,, same for the mants

OpenStudy (danjs):

(1/2)*(1/2)

OpenStudy (anonymous):

white and tan / black and blue = 2/4=1/2?

OpenStudy (anonymous):

is that right??

OpenStudy (anonymous):

because that is what i think?

OpenStudy (anonymous):

@DanJS ??

OpenStudy (danjs):

no, He picks a shirt, there is a 1/2 chance he picks the right color... If he happens to pick the right one, then there is a 1/2 chance he picks the right pants

OpenStudy (anonymous):

so then the answer would be...??

OpenStudy (danjs):

a half times a half or one fourth

OpenStudy (anonymous):

and that would be my answer?

OpenStudy (anonymous):

@DanJS ??

OpenStudy (danjs):

yes, he has a 1 in 4 chance of picking both things correctly